maximum number of atoms.16 g of N216 g of co216 g of CH416 g of O2

Grade:12th pass

1 Answers

Arun
6 years ago
First find the mole of each
N2 = 16/28 = 0.5714
CO2 = 16/44 = 0.3636
CH4 = 16/16 = 1
O2 = 16/3W = 0.5
Since mol of Ch4 is maximum, hence there will be maximum no. of atoms in case of CH4